OpenStudy (anonymous):

Find the logarithm log{2/3] (8/27)

OpenStudy (anonymous):

\[\log_{2/3} 8/27\] Log to the base 2/3 of 8/27

OpenStudy (anonymous):

So we can re-write this as an exponent\[2/3^x = 8/27\]. Or\[2^x/3^x\] So, this question is asking log(2/3) (8/27), or what power do we have to raise 2/3 by to get 8/27. Conveniently, 2^3 =8, and 3^3=27, so (2/3)^3 = 8/27. So the log(2/3) (8/27) =3 Hope this helps.

OpenStudy (anonymous):

Sorry, couldn't get that equation writing to work.

Directrix (directrix):

Let log{2/3] (8/27) = x. By definition of logarithms, then (2/3) ^ x = 8/27 @Ratedover Do you agree so far?

Directrix (directrix):

(2/3) ^ x = 8/27 (2/3) ^ x = ( 2 / 3 ) ^ 3 x = ? @Ratedover